Drugtech has produced another new drug, a vaccine for the common cold! It has received a...

  1. Drugtech has produced another new drug, a vaccine for the common cold! It has received a patent for the drug’s formula and thus can operate as a monopolist until the patent expires. Drugtech faces demand of P=120-5Q for the vaccine and produce each shot of the vaccine at a constant marginal cost of 10 dollars. Hint: MR=120-10Q. Show ALL work.
  1. What is the optimal quantity it should produce?
  2. What is the optimal price it should charge?
  3. Graph the market for the vaccine, labeling price, quantity, and calculate areas of consumer & producer surplus.
Answer #1


HP: 120 -10 Mc = 10 a) At equiliboiu MCMR 120-109 210 -) (@=ll units b) b = 120-50 = 120-5(11) = $65 consumes sueples A B C y